# German Research Center for Artificial Intelligence, Kaiserslautern, Germany

### v1

@@ 1 Introduction
Multi-sensor Machine Learning (ML) models for Earth Observation (EO) use data from diverse sources like satellites and ground-based sensors, providing a comprehensive view of the Earth. Recently, the literature has focused on using these multi-sensor models to improve accuracy with respect to single-sensor models [10]. However, when including more data sources, there is a greater chance that one of the sources will face technical issues and miss some data.
Missing data is inherent in the EO domain, hindering accurate predictions and introducing biases. Moreover, ML models, even adaptive models such as transformers, are not naturally robust to missing data [8]. In Satellite Image Time Series (SITS), cloudy conditions hinder the availability of optical images, negatively affecting model predictions [12]. The drop in performance has also been observed when all the features of a specific sensor are unavailable in land-use classification [6] and vegetation applications [9].
The design of ML models aligned to the missing data scenario has been vaguely explored in the literature. Garnot et al. [12] present a Temporal Dropout (TempD) technique that randomly drops some time-steps in SITS data, obtaining a more robust model to missing time-steps. However, the TempD is focused on single-sensor input [12]. On the other hand, Francis et al. [4] proposed a sensor invariant model that can yield predictions with similar accuracy, regardless of the sensor given as input. However, the proposed model focuses on static images [4]. Motivated by these works, we introduce two prediction methods tailored to the multi-sensor time-series scenario. Our main research questions are: how to handle missing sensors in models at prediction time? how to increase the prediction robustness of model to missing sensors? We use three multi-sensor time series EO datasets to validate our methods and show that we can increase the model robustness compared to previously proposed approaches.
@@ 2 Handling missing data in multi-sensor learning
Multi-sensor learning considers the case when multiple set of features, coming from different sensors, are used as input data in predictive ML models. The main objective is to corroborate the information on individual sensor observations and improve the predictive performance [10]. The literature in EO evidenced that using a multi-sensor perspective is crucial to enrich the input data and improve model performance [6, 12]. However, an appropriate modeling is necessary to get the most advantage of the multi-sensor data [8, 10].
When multi-sensor ML models are deployed in real environments, the assumption that sensor data is persistently available cannot be taken. During inference, the occurrence of missing sensors leads to a scenario for which the model is unprepared. Since re-training can be impractical due to limited resources, some techniques can help to handle in a better way the missing sensor data scenario. In the following, we describe such techniques commonly used in the literature.
@@
A more sophisticated way to fill in the missing sensors is to learn a function that predicts all the missing sensor features. Such techniques include cross-modal auto-encoders and translation models [14]. The latter method has been applied to EO data by translating a radar to optical image with cycleGAN [3], or using multiple optical sensors at different times to reconstruct the optical features at a specific time [13]. Here, the practitioner has to define the entire model that reconstruct, as well as its optimization.
@@ 3.1 Input Sensor Dropout (ISensD)
Figure 1: Illustration of the ISensD method with three sensors as input. As the SensD technique is applied batch-wise, we present three batch examples and the associated randomly masked sensors.
Inspired by recent works that use dropout as data augmentation [12, 18], we introduce a Sensor Dropout (SensD) technique at the input-level. The idea is to randomly mask out all input features of some sensors during training. Consider the set of sensor features (input data) for a sample $i$ as $\mathcal{X}^{(i)}=\{\mbox{{X\/}}^{(i)}_{s}\}_{s\in\mathcal{S}}$, with $\mathcal{S}$ the set of all available sensors. Then, the masked input features can be expressed by $\tilde{\mathcal{X}}^{(i)}=\{d_{s}^{(i)}\cdot\mbox{{X\/}}_{s}^{(i)}\}_{s\in\mathcal{S}}$, with $d_{s}^{(i)}\sim\text{Bernoulli}(r)$ and $r\in[0,1]$ the SensD ratio that controls how much the sensor features will be masked out. In addition, we introduce a special case of the previous technique that does not rely on the hyperparameter $r$. The alternative masking technique lists all the possible combinations of missing sensors ($12^{|\mathcal{S}|}-$ 1) and then selects one randomly. We apply these techniques in a batch-wise manner, with a different random masking $d_{s}^{(i)}$ at each batch. An illustration of the masking is shown in Fig. 1. Finally, the masked features are given as input to a multi-sensor model implementing input-level fusion [10], i.e. $\hat{y}^{(i)}={\mbox{{M}}}_{\theta}({\mbox{{concat}}}(\tilde{\mathcal{X}}^{(i)}))$, with ${\mbox{{M}}}_{\theta}$ the model function parameterized by $\theta$.
@@ 3.2 Ensemble Sensor Invariant (ESensI)
Inspired by recent works that use sensor invariant models in the context of EO [4], we propose a model with sensor invariant layers. The idea is to use an ensemble multi-sensor model with a sensor invariant prediction head. Consider the standard ensemble-based model with sensor-dedicated encoders $z_{s}^{(i)}={\mbox{{E}}}_{\theta_{s}}(\mbox{{X\/}}_{s}^{(i)})\in\mathbb{R}^{d}$, and sensor-dedicated prediction heads $\hat{y}_{s}^{(i)}={\mbox{{P}}}_{\phi_{s}}(z_{s}^{(i)})$. Then, the parameters of each sensor-dedicated model $\Theta_{s}=\{\theta_{s},\phi_{s}\}$ are optimized by a loss function $\mathcal{L}$
@@
Since these sensor-dedicated parameters $\Theta_{s}$ are optimized independently of each other, we incorporate the notion of multi-sensor information by using a shared prediction head in all sensors: $\hat{y}_{s}^{(i)}={\mbox{{P}}}_{\phi}(z_{s}^{(i)}+\rho_{s})\ \forall s\in\mathcal{S}$. We use a learnable sensor encoding vector $\rho_{s}\in\mathbb{R}^{d}$ inspired by the channel group encoding in [16]. The purpose of the sensor encoding vector is to give the model a notion of awareness of which sensor features are being fed for prediction. An illustration of this method is shown in Fig. 2. Since the features $z_{s}$ can have quite different magnitudes, we consider the case in which ${\mbox{{P}}}_{\phi}$ incorporates a normalization in its first layer. Then, as the parameters of the sensor-dedicated model $\Theta_{s}=\{\theta_{s},\rho_{s},\phi\}$ are attached by $\phi$ (the shared prediction head), these are jointly optimized as
@@
Figure 2: Illustration of the ESensI method with three sensors as input. The main component of the method is the shared prediction head, which makes the ensemble-based aggregation a sensor-invariant ML model.
@@ 4.1 Datasets
CropHarvest (CH)
We use the cropharvest dataset for multi-sensor crop recognition [17], in a multi-crop version [9]. This involves a classification task in which the crop-type (between 10 crop-type groups including non-crop) at a given location during a particular season is predicted. The dataset has 29,642 samples around the globe between 2016 and 2022. The temporal sensors are multi-spectral optical (from Sentinel-2), radar (from Sentinel-1), and weather (from ERA5). These were re-sampled monthly during the crop growing year. An additional static sensor is the topographic information. All sensor features were spatially interpolated to a pixel resolution of 10 m.
Live Fuel Moisture Content (LFMC)
We use a dataset for multi-sensor moisture content estimation [11]. This involves a regression task in which the vegetation water (moisture) per dry biomass (in percentage) in a given location at a specific time is predicted. There are 2,578 samples from the western US collected between 2015 and 2019. The temporal sensors are multi-spectral optical (from Landsat-8) and radar (from Sentinel-1). These features were re-sampled monthly during a four-month window prior to the moisture measurement. Additional static sensors are the topographic information, soil properties, canopy height, and land-cover class. All features were interpolated to a pixel resolution of 250 m.
Particulate Matter 2.5 (PM25)
We use a dataset for multi-sensor PM2.5 estimation [2]. This involves a regression task in which the concentration of PM2.5 (particles measure 2.5 microns or less in diameter) in the air (in $ug/m^{3}$) in a particular city at a specific moment is predicted. There are 167,309 samples in five Chinese cities between 2010 and 2015. The temporal sensor features are atmospheric conditions, atmospheric dynamics, and precipitation. These sensors are captured at hourly resolution. We consider a three-day window of information before the day of the PM2.5 measurement.
@@ 4.2 Baselines
For the input-level fusion methods, we align the features of the temporal and static sensors by repeating the static features across the time series, as commonly done in EO [17, 12].
@@ 4.3 Experimental setting
We assess the methods with a 10-fold cross-validation by simulating missing sensors in the validation fold. The assessment consists of having fewer sensors at validation time than during training to make the prediction. We experiment with different degrees of missingness controlled by a percentage of samples that are affected by missing sensors. We consider that sensors with temporal features might be missing, leaving the static sensors as permanently available information. For evaluating model robustness, we use the Performance Robustness Score (PRS) proposed in [5]. This metric is based on the predictive error with missing sensors relative to the predictive error when using all sensors:
@@
For the predictive performance, we use the F1 Macro (F1) macro in classification, and the Coefficient of Determination ($R^{2}$) in regression tasks. All metrics are upper bounded by 1, and a higher value means a better result.
@@ 4.4 Results
We only report the results of the best dropout ratio in the SensD, which is 20% in CH data, 80% in the LFMC data, and 80% in the PM25 data. For the ESensI we include the sensor encoding vector in CH data, in LFMC we add the sensor encoding with normalized vectors, and in PM25 the vector is not included. All the results in the CH data are in Sec. 4.5, while the rest is in the appendix.
@@
In Fig 3, 4, and 5 we display the PRS of the compared methods in the CH, LFMC, and PM25 data respectively. Overall, we observe that Ensemble and ESensI methods are the most robust when increasing the level of missing sensors in the datasets. In addition, Input and ITempD methods have a linear decrease in the robustness with different slopes, depending on the dataset and which sensor is missing. For instance, in the CH and LFMC data, Fig 3 and 4, the decrease is higher when the optical sensor is missing. Besides, in the PM25 data, these methods have an exponential decrease in the robustness.
The PRS results highlight that, by increasing the percentage of missing sensor data, our proposals show a stronger robustness (PRS decreases by a smaller amount) in all cases. The ISensD robustness is significantly increased regarding its direct baselines, i.e. Input and ITempD methods. However, the best robustness is achieved by the Ensemble method, which in some cases (Fig. 5) is improved by the ESensI method. This illustrates that the proposed methods can increase the robustness of multi-sensor models. The same pattern is observed in the predictive performance results (with F1 and $R^{2}$), presented in the appendix.
Table 1: Predictive performance on a full-sensor evaluation, i.e. no missing features, on the different datasets. F1 in classification and $R^{2}$ in regression.
@@
LFMC $0.714$ $\mathbf{0.717}$ $0.650$ $0.313$ $0.540$ $0.553$ $0.326$
PM25 $\mathbf{0.917}$ $0.882$ $-0.354$ $0.308$ $0.512$ $0.471$ $0.232$
In general, the PRS results indicate that the Ensemble method is quite robust without the sensor invariant component. However, when we analyze the predictive performance of the methods with all sensors available, as shown in Table 1, it illustrates that the Ensemble method is among the worst in performance, with the Input method at the top. In this table, it can be seen that the ESensI method improves in two out of three datasets in comparison to the Ensemble method. However, the ISensD methods obtain a fairly poor performance, reflecting the limitation of the SensD technique when all sensors are available.
@@ 4.5 Ablation study
We assess different method configurations in the CH data. Fig. 6 displays the predictive performance when different values of the SensD ratio are used in the ISensD method. There is no clear pattern in the results when this parameter increases or decreases. However, the F1 scores with different SensD ratios oscillate around the “no ratio” version, showing its capacity as a non-parameterizable alternative. For the ESensI method, in Table 2 we show the predictive performance effect of different model components. In this case, it reflects that simply sharing weights is not enough, but rather the other components must be added.
Figure 6: F1 scores of the ISensD method with different SensD ratio values in the CH data. The dashed lines show the no ratio version of the ISensD method.
Table 2: F1 results of different configurations in the ESensI model for the CH data. The results are obtained in the full-sensor evaluation.
Shared weights Sensor encoding Normalization Others F1 score
- - - $0.603$
$\checkmark$ - - $0.595$
$\checkmark$ $\checkmark$ - $0.606$
$\checkmark$ $\checkmark$ $\checkmark$ addition $0.598$
$\checkmark$ $\checkmark$ $\checkmark$ concatenation $0.595$

### v2

@@ 1 Introduction
Multi-sensor Machine Learning (ML) models for Earth Observation (EO) use data from diverse sources, providing a comprehensive view of the Earth. These sources can be from ground-based to satellite-based sensors. Recently, the literature has focused on using these multi-sensor models to improve accuracy regarding single-sensor models [17]. However, by including more sensors, there is a greater chance that one can face technical issues and miss some data [13].
Missing data is inherent in the EO domain, hindering accurate predictions and introducing biases. Moreover, ML models, even adaptive models such as transformers, are not naturally robust to missing data [14]. In Satellite Image Time Series (SITS), cloudy conditions hinder the availability of optical images, negatively affecting model predictions [20]. The decline in performance has also been observed when all features of a specific sensor are unavailable in land-use classification [11] and vegetation applications [15].
The design of ML models aligned to the missing data scenario has been vaguely explored in the literature. Garnot et al. [20] present a Temporal Dropout (TempD) technique that randomly drops some time-steps in SITS data, obtaining a more robust model to missing time-steps. However, the TempD is focused on single-sensor input [20]. On the other hand, Francis et al. [7] proposed a sensor invariant model that can yield predictions with similar accuracy, regardless of the sensor given as input. However, the proposed model focuses on just static images. Motivated by these works, we introduce two prediction methods tailored to the multi-sensor time-series scenario.
The main research questions of our work are: how to handle missing sensors in models at prediction time? how to increase the prediction robustness of model to missing sensors? We use three multi-sensor time series EO datasets to validate our methods in scenarios of missing sensor data. In addition, we show that we can increase the model robustness compared to four methods from the literature.
@@ 2 Handling missing data in multi-sensor learning
Multi-sensor learning considers the case when multiple set of features (from different sensors) are used as input in predictive ML models. The main objective is to corroborate the information on individual observations [17]. The EO literature evidence that using a multi-sensor perspective is crucial to enrich the input data and improve model performance [11, 20, 16]. However, an appropriate modeling is necessary to get the most advantage of the multi-sensor data [14, 18, 17].
When multi-sensor ML models are deployed in real environments, the assumption that sensor data is persistently available cannot be taken. During inference, the occurrence of missing sensors leads to a scenario for which the model is unprepared. Since re-training can be impractical due to limited resources, some techniques can help to handle in a better way the missing of sensor data. In the following, we describe such techniques commonly used in the literature.
@@
A more sophisticated way to fill in the missing sensors is to learn a function that predicts all the missing sensor features. Such techniques include cross-modal auto-encoders and translation models [22]. The latter method has been applied to EO data by translating a radar to optical image with cycleGAN [5], or using multiple optical sensors at different times to reconstruct the optical features at a specific time [21]. Here, the practitioner has to define the entire reconstruction model, as well as its optimization.
@@ 3.1 Input Sensor Dropout (ISensD)
Figure 1: Illustration of the ISensD method. We show three batch examples and the associated randomly masked sensors based on the SensD technique.
Inspired by recent works that use dropout as data augmentation [20], we introduce a Sensor Dropout (SensD) technique at the input-level. The idea is to randomly mask out all input features of some sensors during training. Consider the set of sensor features (input data) for a sample $i$ as $\mathcal{X}^{(i)}=\{\mbox{{X\/}}^{(i)}_{s}\}_{s\in\mathcal{S}}$, with $\mathcal{S}$ the set of all available sensors. Then, the masked input features can be expressed by $\tilde{\mathcal{X}}^{(i)}=\{d_{s}^{(i)}\cdot\mbox{{X\/}}_{s}^{(i)}\}_{s\in\mathcal{S}}$, with $d_{s}^{(i)}\sim\text{Bernoulli}(r)$ and $r\in[0,1]$ the SensD ratio. This ratio that controls how much the sensor features will be masked out. In addition, we introduce a special case of the previous that does not rely on the hyperparameter $r$. The alternative masking technique lists all the possible combinations of missing sensors ($12^{|\mathcal{S}|}-$ 1) and then selects one randomly, which we called No Ratio (NR). We apply these techniques in a batch-wise manner, with a different random masking $d_{s}^{(i)}$ at each batch. An illustration of the masking is shown in Fig. 1. Finally, the masked features are given as input to a multi-sensor model implementing input-level fusion [17], i.e. $\hat{y}^{(i)}={\mbox{{M}}}_{\theta}({\mbox{{concat}}}(\tilde{\mathcal{X}}^{(i)}))$, with ${\mbox{{M}}}_{\theta}$ the model function parameterized by $\theta$. In Sec. 4 we present the model used in ${\mbox{{M}}}_{\theta}$. Finally, the parameters of the model are optimized regarding the ground truth $y^{(i)}$ in the training set by
$\hat{\theta}=\arg\min_{\theta}\sum_{i=1}^{N}{\mbox{{L}}}(y^{(i)},\hat{y}^{(i)})\ .$ (1)
@@ 3.2 Ensemble Sensor Invariant (ESensI)
Inspired by recent works that use sensor invariant models in the context of EO [7], we propose a model with sensor invariant layers. The idea is to use an ensemble multi-sensor model with a sensor invariant prediction head. Consider the standard ensemble-based model with sensor-dedicated encoders $\vec{z}_{s}^{(i)}={\mbox{{E}}}_{\theta_{s}}(\mbox{{X\/}}_{s}^{(i)})\in\mathbb{R}^{d}$, and sensor-dedicated prediction heads $\hat{y}_{s}^{(i)}={\mbox{{P}}}_{\phi_{s}}(\vec{z}_{s}^{(i)})$. Then, the parameters of each sensor-dedicated model $\Theta_{s}=\{\theta_{s},\phi_{s}\}$ are optimized by a loss function $\mathcal{L}$
@@
Since these sensor-dedicated parameters $\Theta_{s}$ are optimized independently of each other, we incorporate the notion of multi-sensor information by using a shared prediction head in all sensors: $\hat{y}_{s}^{(i)}={\mbox{{P}}}_{\phi}(\vec{z}_{s}^{(i)}+\vec{\rho}_{s})\ \forall s\in\mathcal{S}$. We use a learnable sensor encoding vector $\vec{\rho}_{s}\in\mathbb{R}^{d}$ inspired by the channel group encoding in [24]. The purpose of the sensor encoding vector is to give the model a notion of awareness of which sensor features are being fed for prediction. An illustration of this method is shown in Fig. 2. Since the features $z_{s}$ can have quite different magnitudes, we consider the case in which ${\mbox{{P}}}_{\phi}$ incorporates a normalization in its first layer. Then, as the parameters of the sensor-dedicated model $\Theta_{s}=\{\theta_{s},\rho_{s},\phi\}$ are attached by $\phi$ (the shared prediction head), these are jointly optimized as
@@
Figure 2: Illustration of the ESensI method. The main component is the shared prediction head, which makes the ensemble-based a sensor-invariant model.
@@ 4.1 Datasets
CropHarvest (CH).
We use the cropharvest dataset for multi-sensor crop recognition [25], in a multi-crop version [15]. This involves a classification task in which the crop-type (between 10 crop-type groups including non-crop) at a given location during a particular season is predicted. The dataset has 29,642 samples around the globe between 2016 and 2022. The temporal sensors are multi-spectral optical (from Sentinel-2), radar (from Sentinel-1), and weather (from ERA5). These were re-sampled monthly during the crop growing year. An additional static sensor is the topographic information (from SRTM’s DEM). All sensor features are spatially interpolated to a pixel resolution of 10 m.
Live Fuel Moisture Content (LFMC).
We use a dataset for multi-sensor moisture content estimation [19]. This involves a regression task in which the vegetation water (moisture) per dry biomass (in percentage) in a given location at a specific time is predicted. There are 2,578 samples from the western US collected between 2015 and 2019. The temporal sensors are multi-spectral optical (from Landsat-8) and radar (from Sentinel-1). These features were re-sampled monthly during a four-month window before the moisture measurement. Additional static sensors are the topographic information (from NED), soil properties (from Unified North American Soil Map), canopy height (from GLAS’s lidar), and land-cover class (from GlobCover). All features are interpolated to a pixel resolution of 250 m.
Particulate Matter 2.5 (PM25).
We use a dataset for multi-sensor air quality estimation [3]. This involves a regression task in which the concentration (in $ug/m^{3}$) of PM2.5 in the air (particles that are 2.5 microns or less in diameter) in a city at a specific moment is predicted. There are 167,309 samples in five Chinese cities between 2010 and 2015. The temporal sensor features are atmospheric conditions, atmospheric dynamics, and precipitation. These sensors are captured at hourly resolution. We consider a three-day window of information before the day of the PM2.5 measurement.
@@ 4.2 Baselines
For the input-level fusion methods, we align the features of the temporal and static sensors by repeating the static features across the time series, as commonly done in EO [25, 20, 18]. Besides, for the model ${\mbox{{M}}}_{\theta}$, and ${\mbox{{E}}}_{\theta}$ for temporal features (Sec. 3) we use a 1D CNN. We show results with alternative model architectures in the appendix. For the prediction head ${\mbox{{P}}}_{\phi}$, and static features (in Feature and Ensemble), we use an MLP encoder. Two layers with 128 units and a 20% of dropout are used. We use the ADAM optimizer with a batch-size of 128 over the cross-entropy in classification and mean squared error in regression.
@@ 4.3 Experimental setting
We assess the methods with a 10-fold cross-validation by simulating missing sensors in the validation fold. The assessment consists of having fewer sensors at validation time than during training to make the prediction. We experiment with different degrees of missingness controlled by a percentage of samples that are affected by missing sensors. We consider that sensors with temporal features might be missing, leaving the static sensors as permanently available information. For evaluating the model robustness, we use the Performance Robustness Score (PRS) proposed in [10]. This metric is based on the predictive error with missing sensors relative to the predictive error when using all sensors:
@@
For the predictive performance, we use the F1 Macro (F1) macro in classification and the Coefficient of Determination ($R^{2}$) in regression tasks.
@@ 4.4 Results
We report the results of the best dropout ratio in the SensD, which is 60% in CH data, 25% in the LFMC data, and 45% in the PM25 data. For the ESensI we include the sensor encoding vector in CH data, in LFMC we add the sensor encoding with normalized vectors, and in PM25 the vector is not included. Ablation studies on these parameters are shown in Sec. 4.5.
@@
In Fig 3, 4, and 5 we display the PRS of the compared methods in the CH, LFMC, and PM25 data respectively. Overall, we observe that Ensemble and ESensI methods are the most robust when increasing the level of missing sensors in the datasets. In addition, Input and ITempD methods have a linear decrease in the robustness with different slopes, depending on the dataset and which sensor is missing. For instance, in the CH and LFMC data, Fig 3 and 4, the decrease is higher when the optical sensor is missing. Besides, in the PM25 data, these methods have an exponential decrease in robustness.
The PRS results highlight that, by increasing the percentage of missing sensor data, our proposals show stronger robustness (PRS decreases by a smaller amount) in all cases. The ISensD robustness is significantly increased regarding its direct baselines, i.e. Input and ITempD methods. However, the best robustness is achieved by the Ensemble method, which in some cases (Fig. 5) is improved by the ESensI method. This illustrates that the proposed methods can increase the robustness of multi-sensor models. The same pattern is observed in the predictive performance results (with F1 and $R^{2}$), presented in the appendix.
Table 1: Predictive performance on a full-sensor evaluation, i.e. no missing features, on different datasets. The F1 is shown in classification and $R^{2}$ in regression.
@@
LFMC $0.714$ $\mathbf{0.717}$ $0.650$ $0.313$ $0.655$ $0.545$ $0.326$
PM25 $\mathbf{0.917}$ $0.882$ $-0.354$ $0.308$ $0.525$ $0.496$ $0.232$
In general, the PRS results indicate that the Ensemble method is quite robust without the sensor invariant component, i.e. ESensI. However, when we analyze the predictive performance with all sensors available, in Table 1, it reflects that the Ensemble method is among the worst in predictive performance, with the Input method at the top. Here, we observe that the ESensI method improves in two out of three datasets in comparison to the Ensemble method. However, the ISensD method performs worse than its counterpart, Input, reflecting the limitation of the SensD technique when all sensors are available. We suspect that the model is focusing more on the missing data, ignoring the full-sensor scenario, as this occurs only once within all the randomized scenarios.
@@ 4.5 Ablation study
We assess different configurations in the methods presented. Fig. 6 displays the predictive performance when different values of the SensD ratio are used in the ISensD method. Depending on the missing sensor, increasing the ratio increases or decreases the performance, while for the full-sensor scenario, it always tends to decrease. However, the results with different SensD ratios variate around the NR version, showing its good capacity as a parameter-free alternative. For the ESensI method, we show the predictive performance with different model components in Table 2. In this case, it reflects that simply sharing weights is not optimal in the CH and LFMC data, but rather additional components should be added. On the contrary, in PM25, it is better to use shared weights in the model prediction head.
Figure 6: Predictive performance results of the ISensD method with different SensD ratio values. The dashed lines show the no ratio version of the ISensD method.
Table 2: Predictive performance results of different configurations in the ESensI model. The results are obtained in the full-sensor evaluation.
Shared weights Sensor encoding Normalize Others CH           (F1) LFMC           ($R^{2}$) PM25           ($R^{2}$)
- - - $0.603$ $0.313$ 0.3080.308{0.308}0.308
$\checkmark$ - - $0.595$ $0.281$ $\mathbf{0.281}$
$\checkmark$ $\checkmark$ - $\mathbf{0.606}$ 0.3090.309{0.309}0.309 $0.274$
$\checkmark$ $\checkmark$ $\checkmark$ add $0.598$ $\mathbf{0.317}$ $0.241$
$\checkmark$ $\checkmark$ $\checkmark$ concat $0.595$ $0.297$ $0.196$
5 Related work
In the following, we discuss some related works in the multi-sensor EO literature.
Multi-sensor learning.
Recent research in self-supervised learning and foundational models has placed the usage of multi-sensor data as a basis. Most of these works use feature-level fusion with sensor-dedicated encoders to handle sensors with different spatial resolutions. For instance, in Chen et al. [4], a pure Transformer model is introduced that uses cross-modal contrastive and reconstruction learning as pre-training. While in OmniSat [2], a similar model is introduced based on contrastive and reconstruction learning but adapted to sensors with different temporal resolutions. In Guo et al. [8], Skysense is proposed as a geospatial foundational model through cross-modal contrastive learning and knowledge distillation (via a student-teacher framework). Additionally, some works compare different sensor fusion strategies for pixel-wise [16] and image-wise [6] crop classification with SITS. In any of these models, the SensD technique can be integrated by just randomly masking out all input features from the multi-sensor data. While for the sensor-invariant component, it can be integrated in any of the sensor-dedicated layers by just sharing some parameters.
Random masking.
Masking out the input data has been used in the literature as an augmentation technique during training. For instance, by randomly dropping time-steps of SITS [20], features in random modalities [24], or patches in random modalities [2, 4]. In OmniSat [2], a masking vector is learned that replaces the masked data, while in Presto [24] and Chen et al. [4] the masked data is ignored by using masked attention models. In these works, Transformer models are used with a feature-level fusion strategy.
Sensor invariant.
The introduced ESensI is directly motivated by Francis [7], which uses the same layers for different sensor descriptors, including a block that generates all possible sensor permutations (permutation block). However, in that work, they just focus on satellite image cloud segmentation. In Hackstein et al. [9], different ways of using shared layers (common blocks) are compared for cross-sensor reconstruction. The results show that shared layers are more effective in the decoder than the encoder, as we used. Besides, shared layers allow reducing the computational resources of the multi-sensor model.